OpenStudy (anonymous):

How do you simplify the square root of -300?

OpenStudy (anonymous):

If you are going to simplify it, you want to split it into 2 square roots. In this case, it will be sqrt of -3 multiplied by sqrt of 100. Simplifying will get you 10sqrt(-3).

OpenStudy (mathmate):

\( \sqrt{-300} = \sqrt{2^2*5^2*3*(-1)} = 2*5*\sqrt{3} \sqrt{-1}=10\sqrt{3}i\)

OpenStudy (anonymous):

Should the i be on the outside?

OpenStudy (mathmate):

Yes, the i should be, and is on the outside. Unfortunately it is not very clear.

OpenStudy (anonymous):

the options for the answers are \[3i \sqrt{10}\] or \[10i \sqrt{3}\]

OpenStudy (mathmate):

Which one is equivalent to what Veillant or I gave above?

OpenStudy (anonymous):

I have no idea, what do you think?

OpenStudy (anonymous):

thanksssss !

OpenStudy (mathmate):

Does \( 10i\sqrt{3} = 10\sqrt{3}i \) or \( 3i\sqrt{10}=10\sqrt{3}i \) ?

OpenStudy (anonymous):

not sureeee

OpenStudy (anonymous):

I've always been told when you root a negative, always take the i out. The first thing that came to me to simplify was sqrt3 and sqrt100. If anything, just leave 10isqrt3.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

No problem!